OpenStudy (anonymous):

Give the exact solutions. y^2 -16y+64=16 y=

OpenStudy (anonymous):

\[y^2 -16y+64=16\]

OpenStudy (anonymous):

(y-8)^2 = 16 sqrt both sides y-8 = +4 or -4 y = +12 or +4